OpenStudy (maddy1251):

What would log(10^3x^7)= ? Would it be 3+7log(x)?

OpenStudy (maddy1251):

\[\log(10^{3}x^{7})\]

Nnesha (nnesha):

no first apply product rule and then power rule

Nnesha (nnesha):

product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition <----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\]

OpenStudy (maddy1251):

\[so 3\log(10)+7\log(x) \]

OpenStudy (maddy1251):

Oh

Nnesha (nnesha):

yes looks good

OpenStudy (maddy1251):

Would that then be \[21\log(x)\]

Nnesha (nnesha):

no hmm you can't multiply 3 and 7

OpenStudy (maddy1251):

See none of these look right to me, unless I am forgetting something.

Nnesha (nnesha):

huh hmm

OpenStudy (maddy1251):

3+7log(x) was the only one I thought could be possible..Even thought I knew it wasn't right, not like what we wrote above.

OpenStudy (anonymous):

log10= 1

OpenStudy (maddy1251):

@ArifL so the 3 log(10) would be 3 (1)?

OpenStudy (anonymous):

yup

OpenStudy (maddy1251):

1+7log(x)?

OpenStudy (anonymous):

what is 3x1

OpenStudy (maddy1251):

oh 3.. haha 3+7log(x)

OpenStudy (maddy1251):

@ArifL Thank you! :D

OpenStudy (anonymous):

welcome